OpenStudy (kittiwitti1):

226Ra has a half-life of 1599 years. How much is left after 1000 years if the initial amount was 10 g?

OpenStudy (anonymous):

i dont know sorry

OpenStudy (kittiwitti1):

\[y=ae^{bx}\]\[x=1599\]\[a=10g\]\[y=?\]

OpenStudy (kittiwitti1):

Meh @anabelle

OpenStudy (anonymous):

sos

OpenStudy (kittiwitti1):

lol

OpenStudy (anonymous):

XD

OpenStudy (kittiwitti1):

I did it like this: (0,10)\[0=10e^{b \times1599}\]\[0=e^{1599b}\]

OpenStudy (kittiwitti1):

... aaaand now I'm stuck

OpenStudy (unklerhaukus):

\[[\text{Ra}](t)=[\text{Ra}]_0 (\tfrac12)^{t/t_{1/2}}\]

OpenStudy (kittiwitti1):

Okay?

OpenStudy (unklerhaukus):

you have the half life so you should the half life formula not exponential decay formula the difference is that the base is 1/2 not e

OpenStudy (unklerhaukus):

t is the time (you can use x if you really want to) t_1/2 is the half life [Ra](t) is the concentration of Radon at time t [Ra]_0 is the initial concentration of Radon

OpenStudy (kittiwitti1):

o.o We did it like this in class:\[y=ae^{bx}\]\[\frac{1}{2}=ae^{b\times 0}\]\[a=\frac{1}{2}\]\[y=\frac{1}{2}e^{bx}\]Is that what you were referring to?

OpenStudy (kittiwitti1):

If yes, /headdesk I was stupid and didn't realize LOL

OpenStudy (kittiwitti1):

and I meant to type this \[\frac{1}{2}=ae^{bx}\]

OpenStudy (unklerhaukus):

its kinda similar

OpenStudy (unklerhaukus):

\[[\text{Ra}](1000)=10[\text g] (\tfrac12)^{1000/1599}\]

OpenStudy (kittiwitti1):

Uh, can we simplify into easier-to-read variables?

OpenStudy (unklerhaukus):

you can if you wish, but you should have 1/2 as the base not e

OpenStudy (unklerhaukus):

\[\large y(x)=a(\tfrac12)^{x/x_{1/2}}\]

OpenStudy (kittiwitti1):

Ah, okay... what is x 1/2? Is that x (while x is b)?

OpenStudy (anonymous):

just add by tens and if it is enough then you get the answer, from 10 - 100 it is not that hard =)

OpenStudy (unklerhaukus):

t_1/2=x_1/2= the half life

OpenStudy (kittiwitti1):

@carolinagb02 I'm sure these formulas exist for a reason

OpenStudy (kittiwitti1):

@UnkleRhaukus so\[y=a\frac{1}{2}^{bx}\]or am I wrong?

OpenStudy (unklerhaukus):

it depends how you define b

OpenStudy (kittiwitti1):

b is the rate of decay

OpenStudy (unklerhaukus):

i suppose you can have e as the base if \[b=\frac{-\ln(1/2)}{t_{1/2}}=\frac{\ln(2)}{t_{1/2}}\]

OpenStudy (kittiwitti1):

wat

OpenStudy (kittiwitti1):

Maybe I should just go over this in class and get back to you on this lol

OpenStudy (kittiwitti1):

It's 6:45 AM right now... class starts at 7:30

OpenStudy (kittiwitti1):

I still have to sleep T^T

OpenStudy (unklerhaukus):

if you really wanted to use that first formula \[y=ae^{bx}\] you find \(b\) by \(b=\ln2/\text{half life}\)

OpenStudy (kittiwitti1):

I have to go, but I'll keep that in mind

OpenStudy (unklerhaukus):

x is 1000

OpenStudy (kittiwitti1):

Actually, we do it like this: \[y=ae^{bx}\]\[1=2e^{1599(x)}\]\[\frac{1}{2}=e^{1599x}\]\[ln\frac{1}{2}=1599x\]\[x\approx. -4.3349\]

OpenStudy (unklerhaukus):

\[\large y(t)=y_0e^{-\ln 2 \frac{t}{t_{1/2}}}\\ \\ =10e^{-\ln 2\frac{1000}{1599}}\\ \approx 6.5\]

OpenStudy (kittiwitti1):

wat o_o